# On $\nu$ faces of partial alternating sign matrix polytopes

### v1

@@ 5. The $(\nu/\lambda)$-partial ASM polytope is an order polytope and flow polytope
In this section, we show $\mathrm{PASM}(\nu/\lambda,m,n)$ is an order polytope, and so is consequently a flow polytope. We also show how this polytope relates to the ASM-CRY polytope.
@@
We use the corner sum map to construct a map from the polytope $\mathrm{PASM}(\nu/\lambda)$ to the order polytope $\mathcal{O}(P(\nu/\lambda))$, which we show in Theorem 5.4 is an integral equivalence. For $S\subseteq\mathbb{R}$, let $\mathcal{A}(P(\nu/\lambda),S)$ be the set of functions $g:P(\nu/\lambda)\rightarrow S$. In this notation, the order polytope $\mathcal{O}(P(\nu/\lambda))$ is a subset of $\mathcal{A}(P(\nu/\lambda),[0,1])$. Define $\Psi:\mathrm{PASM}(\nu/\lambda)\rightarrow\mathcal{A}(P(\nu/\lambda),\mathbb{R})$ by $a\rightarrow g_{a}$ where $g_{a}(i,j)=c_{ij}$. See Figure 6 for an example.
@@
Figure 6. The map $\Psi$ from a point in the polytope $\mathrm{PASM}(\nu/\lambda)$ to a point in the order polytope $\mathcal{O}(P(\nu/\lambda))$ for $\lambda=(3,1)$ and $\nu=(4,2,2)$. The first step replaces each entry by the northwest corner sum. The second step labels the poset $P(\nu/\lambda)$ with the corresponding values.
@@
The $(\nu/\lambda)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$.
The following lemma shows the map $\Psi$ sends $\mathrm{PASM}(\nu/\lambda)$ into the order polytope. Then in the proof of Theorem 5.4, we show $\Psi$ is an integral equivalence.
@@
Thus, the proof will be complete by showing $c_{ij}\leq c_{i^{\prime}j^{\prime}}$ whenever $i\leq i^{\prime}$ and $j\leq j^{\prime}$. Consider $c_{ij}$ and $c_{i+1,j}$. $c_{i+1,j}=c_{ij}+\displaystyle\sum_{j^{\prime}=1}^{j}X_{i+1,j^{\prime}}$. By (3.1), the partial column sum $\sum_{j^{\prime}=1}^{j}X_{i+1,j^{\prime}}\geq 0$, so $c_{ij}\leq c_{i+1,j}$. The inequality $c_{ij}\leq c_{i,j+1}$ follows analogously. Therefore, $\Psi$ maps $\mathrm{PASM}(\nu/\lambda)$ into the order polytope $\mathcal{O}(P(\nu/\lambda))$. ∎
@@
By definition, the map $\Psi$ is an affine transformation of the form $X\mapsto AX$ where $A$ is a $0,1$-matrix that is lower triangular with ones on the diagonal. So it is a unimodular transformation. By Lemma 5.5, the image of this map is in the order polytope $\mathcal{O}(P(\nu/\lambda))$. Thus, $\Psi$ is a bijection between $\mathrm{PASM}(\nu/\lambda)$ and $\mathcal{O}(P(\nu/\lambda))$ that preserves their respective lattices. This shows that the two polytopes are integrally equivalent. ∎
@@
The normalized volume of $\mathrm{PASM}(\nu/\lambda)$ is $e(P(\nu/\lambda))$ and its Ehrhart polynomial is $L_{\mathrm{PASM}(\nu/\lambda)}(t)=\Omega(P(\nu/\lambda),t+1)$. In particular, $e(P(\nu/\lambda))$ is the number of skew standard Young tableaux, enumerated by the Naruse hook-length formula
@@
We now state a few more definitions from [9] needed for Corollary 5.10, which interprets $\mathrm{PASM}(\nu/\lambda)$ as a flow polytope. The first definition will be used to specify a graph $G_{P(\nu/\lambda)}$ corresponding to the poset $P(\nu/\lambda)$.
@@
The $(\lambda,n)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda)$ is integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$.
@@
By Theorem 5.4, $\mathrm{PASM}(\nu/\lambda)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$. By [9, Theorem 3.14], since the poset $P(\nu/\lambda)$ is strongly planar, $\mathcal{O}(P(\nu/\lambda))$ is integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$. Thus, $\mathrm{PASM}(\nu/\lambda)$ is also integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$. ∎
Our final result describes how $\mathrm{PASM}(\nu/\lambda,m,n)$ relates to the $\lambda$-ASM-CRY polytope $\mathrm{ASMCRY}(\lambda,n)$ of [9] (see Definition 1.2) in the case $m=n$ and $\nu=\delta_{n}:=(n-1,n-2,\ldots,2,1)$.

### v2

@@ 5. The $(\nu/\lambda)$-partial ASM polytope is an order polytope and flow polytope
In this section, we show $\mathrm{PASM}(\nu/\lambda,m,n)$ is an order polytope of a strongly planar graph, and so is consequently a flow polytope. We also show how this polytope relates to the ASM-CRY polytope.
@@
We use the corner sum map to construct a map from the polytope $\mathrm{PASM}(\nu/\lambda,m,n)$ to the order polytope $\mathcal{O}(P(\nu/\lambda))$, which we show in Theorem 5.4 is an integral equivalence. For $S\subseteq\mathbb{R}$, let $\mathcal{A}(P(\nu/\lambda),S)$ be the set of functions $g:P(\nu/\lambda)\rightarrow S$. In this notation, the order polytope $\mathcal{O}(P(\nu/\lambda))$ is a subset of $\mathcal{A}(P(\nu/\lambda),[0,1])$. Define $\Psi:\mathrm{PASM}(\nu/\lambda,m,n)\rightarrow\mathcal{A}(P(\nu/\lambda,m,n),\mathbb{R})$ by $a\rightarrow g_{a}$ where $g_{a}(i,j)=c_{ij}$. See Figure 6 for an example.
@@
Figure 6. The map $\Psi$ from a point in the polytope $\mathrm{PASM}(\nu/\lambda,4,5)$ to a point in the order polytope $\mathcal{O}(P(\nu/\lambda))$ for $\lambda=(3,1)$ and $\nu=(4,2,2)$. The first step replaces each entry by the northwest corner sum. The second step labels the poset $P(\nu/\lambda)$ with the corresponding values.
@@
The $(\nu/\lambda)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda,m,n)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$.
The following lemma shows the map $\Psi$ sends $\mathrm{PASM}(\nu/\lambda,m,n)$ into the order polytope. Then in the proof of Theorem 5.4, we show $\Psi$ is an integral equivalence.
@@
Thus, the proof will be complete by showing $c_{ij}\leq c_{i^{\prime}j^{\prime}}$ whenever $i\leq i^{\prime}$ and $j\leq j^{\prime}$. Consider $c_{ij}$ and $c_{i+1,j}$. $c_{i+1,j}=c_{ij}+\displaystyle\sum_{j^{\prime}=1}^{j}X_{i+1,j^{\prime}}$. By (3.1), the partial column sum $\sum_{j^{\prime}=1}^{j}X_{i+1,j^{\prime}}\geq 0$, so $c_{ij}\leq c_{i+1,j}$. The inequality $c_{ij}\leq c_{i,j+1}$ follows analogously. Therefore, $\Psi$ maps $\mathrm{PASM}(\nu/\lambda,m,n)$ into the order polytope $\mathcal{O}(P(\nu/\lambda))$. ∎
@@
By definition, the map $\Psi$ is an affine transformation of the form $X\mapsto AX$ where $A$ is a $0,1$-matrix that is lower triangular with ones on the diagonal. So it is a unimodular transformation. By Lemma 5.5, the image of this map is in the order polytope $\mathcal{O}(P(\nu/\lambda))$. Thus, $\Psi$ is a bijection between $\mathrm{PASM}(\nu/\lambda,m,n)$ and $\mathcal{O}(P(\nu/\lambda))$ that preserves their respective lattices. This shows that the two polytopes are integrally equivalent. ∎
@@
The normalized volume of $\mathrm{PASM}(\nu/\lambda,m,n)$ is $e(P(\nu/\lambda))$ and its Ehrhart polynomial is $L_{\mathrm{PASM}(\nu/\lambda,m,n)}(t)=\Omega(P(\nu/\lambda),t+1)$. In particular, $e(P(\nu/\lambda))$ is the number of skew standard Young tableaux, enumerated by the Naruse hook-length formula
@@
We now state a few more definitions from [9] that are needed for Corollary 5.10 which interprets $\mathrm{PASM}(\nu/\lambda,m,n)$ as a flow polytope. The first definition will be used to specify a graph $G_{P(\nu/\lambda)}$ corresponding to the poset $P(\nu/\lambda)$.
@@
The $(\lambda,n)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda,m,n)$ is integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$.
@@
By Theorem 5.4, $\mathrm{PASM}(\nu/\lambda,m,n)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$. By [9, Theorem 3.14], since the poset $P(\nu/\lambda)$ is strongly planar, $\mathcal{O}(P(\nu/\lambda))$ is integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$. Thus, $\mathrm{PASM}(\nu/\lambda,m,n)$ is also integrally equivalent to the flow polytope $\mathcal{F}_{G_{P(\nu/\lambda)}}$. ∎
Our final result describes how $\mathrm{PASM}(\nu/\lambda,m,n)$ relates to the $\lambda$-ASMCRY polytope $\mathrm{ASMCRY}(\lambda,n)$ of [9] (see Definition 1.2) in the case $m=n$ and $\nu=\delta_{n}:=(n-1,n-2,\ldots,2,1)$.